please help me to solve this:
find three consecutive integers. the square of the largest one equals the sum of the squares of the other two.

Answer by nadinadan(32)

Answer:
Let's change the words into math operations.
(x+2)^2=(x+1)^2 +x Simplify
x^2+2x+4=x^2+2x+1+x Put all the numbers on one side of the equal sign.
x^2+2x+4-x^2-2x-1-x=0 Simplify
x=3
So the 3 consecutive numbers are 3,4,5.
Verify: 5^2=3^2+4^2 25=9+16